Use the graph to determine whether each system has no solution, one solution, or infinetly many solutions..
1. Y=x-1
Y=-x+1
2. X-y=-4
Y=x+4
3. Y=x+4
2x-2y=2
4. Y=2x-3
2x-2y=2

Answer by Alan3354(69443)

Use the graph to determine whether each system has no solution, one solution, or infinetly many solutions..
1. Y=x-1
Y=-x+1
-----------
These are straight lines, so you need 2 points.
y = x-1
Pick 2 values for x and find y.
Plot the 2 points.
Draw a line thru them.
==============
The intersection (if any) is the solution.
Do the others the same way.
----------------------------------------
2. X-y=-4
Y=x+4
3. Y=x+4
2x-2y=2
4. Y=2x-3
2x-2y=2